J & J Furnace Corp. purchased machinery for $315,000 on May 1, 2017. It is estimated that it will have a useful life of 10 years, salvage value of $15,000, production of 240,000 units, and working hours of 25,000. During 2018, Seceda Corp. uses the machinery for 2,650 hours, and the machinery produces 25,500 units.

From the information given, compute the depreciation charge for 2018 under each of the following methods

  1. Straight-line
  2. Units-of-output
  3. Working hours
  4. Declining-balance
  5. Sum of the years’ digits
Answer #1
Solution:
a. Straight line
Depreciation per year = Cost of an asset - salvage value/useful life of an asset
                                            = $315,000 - $15,000 / 10
                                            = $30,000
Depreciation for the year 2018 = $30,000
b. Units of output
Depreciation base = $315,000 - $15,000 = $300,000
Depreciation charge for 2018 = Depreciation base * Production for the year / Total capacity
                                                         = $300,000*25,500/240,000
                                                         = $31,875
c. Working hours
Depreciation base = $315,000 - $15,000 = $300,000
Depreciation charge for 2018 = Depreciation base * Hours worked for the year / Total capacity in hours
                                                         = $300,000*2,650/25,000
                                                         = $31,800
d. Declining balance
Depreciation rate = (100/10 years * 2 )/100 = 20%  
In 2017 , depreciation = $315,000 *20% *8/12 = $42,000
In 2018 , depreciation = ($315,000 - $42,000 )*20% = $54,600
e. Sum of the years digit
Depreciation base = $315,000 - $15,000 = $300,000
Sum of the year = 1+2+3+4+5+6+7+8+9+10 = 55
Depreciation charge = Remaining useful life if an asset / Sum of years digit * Deprecition base
                                       = 9/55*$300,000
                                       = $49,091
Please upvote , if found the answer useful.
For any clarification, Feel free to reach at comment Section